OpenStudy (anonymous):

Help!! Medal and fan!! Convert the mixed number to an improper fraction. 5 3/20 A.28/20 B.35/20 C.103/20 D.128/20

OpenStudy (anonymous):

the answer is c because 20 * 5=100+3=103

OpenStudy (barreraa):

2 13/20

OpenStudy (anonymous):

103 /20 becomes 5 3/20 plz fan and give me best response